# FADE: A Dataset for Detecting Falling Objects around Buildings in Video

### v1

@@ III-C Dataset division and statistics
Dataset Division. When splitting the dataset, we consider five labels, i.e. video resolution, scene, lighting condition, weather condition, and object category, for each video. Videos in the training set, validation set, and testing set include all of these labels. To better test the generalization performance of the algorithm, the scenes in our training set, validation set, and testing set (except for the scenes captured in rainy days) are non-overlapped.
TABLE II: Area ranges and proportions of the falling objects in our FADE dataset, sorted by ascending order of the area ranges. The video frame size is rescaled to $640\times 480$.
Range of $area$uparrow Number Proportion %
$0^{2}<area\leq 5^{2}$ 37,400 55.34
$5^{2}<area\leq 10^{2}$ 18,882 27.94
$10^{2}<area\leq 15^{2}$ 7,036 10.41
$15^{2}<area\leq 20^{2}$ 2,808 4.15
$area>20^{2}$ 1,459 2.16
Dataset Statistics. We provide some statistical information about the constructed dataset. First, as shown in Figure 5, to make our FADE dataset with abundant falling object categories, we collect the videos covering various falling objects. The percentage of falling objects with different weights is shown in Figure 6. Second, as shown in Table II, most falling objects in our FADE dataset are of small size, where the median area of the falling object is about 20 pixels with the video frame size rescaled to $640\times 480$.
Figure 7: Overview of our proposed method. Where our method is built upon the Faster R-CNN [78]+FPN [79] framework. “C” and “A” denote classifier and anchor regressor, respectively. “Conv” and “AdaConv” indicate conventional convolution and the adaptive convolution [80] layers, respectively.
Figure 8: Details of the moving attention module.
@@
Ethics. We have full ownership of these videos supplied in our FADE dataset, and all of the collected videos have been authorized for free use by the corresponding university. We inform the seven volunteers, who help us drop the objects from buildings, in advance that their personal information (e.g. faces and hands) may appear in the video, and we get their permission. We conduct two improvements to solve the ethical concerns in our dataset. On the one hand, we have checked all the videos and find that there are 32 videos in our dataset have some people appear. The volunteers who help us to collect the videos are captured in 21 of the videos. For the 21 videos, the volunteers are informed of their appearance and have signed on an informed consent form to allow us to use the videos. The link of this signed informed consent form is: https://fadedataset.github.io/FADE.github.io/ethic.html. For the other 11 videos, we remove them from our FADE dataset. On the other hand, we have talked with the corresponding university where we captured the video data and get the authorization. All the videos in our updated FADE dataset can be freely used for the research of falling object detection. We upload the corresponding authorization files to our dataset website. The link is: https://fadedataset.github.io/FADE.github.io/ethic.html. The authorization files of the corresponding university and the volunteers are available at our dataset website https://fadedataset.github.io/FADE.github.io/ethic.html.
@@
Data maintenance and long-term preservation. Our dataset can be downloaded at the website https://fadedataset.github.io/FADE.github.io/download.html. We will upload all the video data of our dataset if the paper could be accepted. Our dataset will be maintained for a long time. The corresponding author will provide the support and can be contacted via e-mail. We plan to add more data to perfect our dataset in future.
@@ III-E Evaluation metrics
We utilize the precision, recall, and F-measure to compare the performance of different methods, including the tasks of MOD, generic object detection, and video object detection. Since the object size in our dataset FADE is small, we treat a detection as a true positive if its IoU with the GT is larger than 0.3. The F-measure is defined as:
@@
In addition, we set the positive real factor $\beta=1$, when we calculate the F-measure.
@@ IV Proposed Method
IV-A Overview
Generic object detection methods are capable of effectively utilizing appearance features to detect objects. However, for the FODB task, the falling objects are usually small and moving at high speed, making it hard to detect them solely based on the appearance features. To address this issue, we propose a method named FADE-Net, which utilizes motion cues to complement appearance features, as well as generate high-quality proposals for small objects. Specifically, we integrate Moving Attention Modules (MAM) with Small-Object Mining RPN (SMRPN) into a generic object detection method, i.e. Faster R-CNN [78]+FPN [79], as shown in Figure 7.
IV-B Moving attention module (MAM)
To begin with, we employ the moving object prediction module which utilizes the MOD method MOG2 [42] to generate moving object masks for consecutive video frames. MOG2 uses a Gaussian mixture model to imitate background for localizing moving objects. A moving object mask provides candidate moving object regions in a frame. The moving object mask is input to several attention modules for generating attention masks. To recall all the possible moving objects, we choose a relative low threshold for MOG2 to produce the moving object masks.
Then, we introduce three attention modules into the backbone as shown in Figure 7. Each attention module takes the feature map from its preceding layer and the moving object mask from the moving object prediction module as input and outputs an attention map which is multiplied with the feature map to highlight the moving object regions in the feature map. The detailed structure of an attention module is shown in Figure 8. Let $F$ be the feature map of size $H\times W\times C$ from the preceding layer of an attention layer. The moving object mask is resized to the size of $H\times W\times 1$, and we denote the resized moving object mask by $M$. We apply max pooling and average pooling to $F$, respectively. The outputs are concatenated with the resized moving object mask and fed to a convolution layer followed by Sigmoid activation to generate a moving attention map. The generation process of the moving attention map (MAP) can be expressed as:
$\begin{split}MAP=\sigma\left(f^{Conv}\left([AvgPool(F),MaxPool(F),M]\right)\right).\end{split}$ (4)
The kernel size of the convolution layer is set to $7\times 7$, and $\sigma$ represents the Sigmoid function.
IV-C Small-Object Mining RPN (SMRPN)
We introduce Small-Object Mining RPN (SMRPN), a region proposal network designed to tackle the limitations of existing RPN approaches [78, 82] in handling small objects. Inspired by Cascade RPN [80], SMRPN performs a multi-stage regression process to refine the proposals iteratively. To address the issue of insufficient and low-quality proposals for small objects, SMRPN employs an area-based anchor mining strategy with a dynamic threshold at the first stage [83]. The threshold can be computed as:
@@ IV-C Small-Object Mining RPN (SMRPN)
$\begin{split}Threshold=\mathrm{max}\left(0.20,0.15+\alpha\cdot\log{\frac{\sqrt{w\cdot h}}{5}}\right)\end{split}$ (5)
where $\alpha$ denotes a scale factor and is set to a default value of 0.2 in our task. The term 5 refers to the minimal area definition of our FADE dataset, ensuring sufficient positive samples for objects of various sizes. The max operation keeps the optimization from being overwhelmed via low quality priors. Moreover, by using the proposed continuous threshold, the model can determine positive samples more smoothly.
Additionally, unlike Cascade RPN, which relies on a single feature pyramid level, SMRPN leverages features from all levels to perform the initial regression, and the information for objects at different scales can be effectively captured accordingly. Following, based on the captured offsets, an adaptive convolution [80] is used to align the anchor and target feature to obtain more refined regression, ultimately generating high-quality proposals for small objects. Consequently, this method is valid for handling the challenges posed by small object detection, paving the way for improving the accuracy of FODB in FADE.
@@ V Experiments
TABLE III: Performance of different modules in the proposed FADE-Net on the testing set of our FADE dataset. The best performances are highlighted in bold.
Faster R-CNN+FPN SMRPN MAM F-measure Precision Recall TRO
$\surd$ 35.52 54.50 26.34 32.44
$\surd$ $\surd$ 53.41 62.67 46.53 39.12
$\surd$ $\surd$ 61.67 59.16 64.40 47.21
$\surd$ $\surd$ $\surd$ 72.03 73.48 70.65 51.75
TABLE IV: Performance of different methods on the testing set of our FADE dataset. The best and second best performances are highlighted in bold and underline, respectively. DLA34 (DCNv2 [84][85] indicates that some of the convolutions in the DLA34 network are replaced by Deformable Convolution v2 (DCNv2). FGFA (FlowNet [86][87], FGFA (PWC-Net [88][87], and FGFA (RAFT [89][87] respectively indicate FGFA based on the optical flow methods of FlowNet, PWC-Net, and RAFT.
Method Type F-measure Precision Recall TRO
FADE-Net (Ours) MOD-based Generic Object Detection 72.03 73.48 70.65 51.75
Faster R-CNN [78]+FPN [79] Generic Object Detection 35.52 54.50 26.34 32.44
YOLOv5 [90] Generic Object Detection 33.64 54.75 24.28 34.09
DLA34 (DCNv2 [84][85] Generic Object Detection 22.57 20.05 25.80 24.66
DETR [91] Generic Object Detection 29.95 49.45 21.48 26.87
swin-B [92] Generic Object Detection 36.93 57.32 27.25 36.58
MOG [93] Moving Object Detection 17.93 12.82 29.82 20.95
MOG2 [42] Moving Object Detection 24.48 18.96 34.53 47.98
GMG [94] Moving Object Detection 2.07 1.14 10.82 14.60
Vibe [8] Moving Object Detection 15.20 14.61 15.85 23.76
CNT [95] Moving Object Detection 12.88 19.17 9.70 17.21
FMOD [96] Moving Object Detection 0.76 6.69 1.36 16.02
KNN [97] Moving Object Detection 0.72 0.36 30.52 34.16
GSOC [98] Moving Object Detection 0.61 0.32 10.15 16.18
LSBP [99] Moving Object Detection 0.12 0.06 6.69 12.36
MEGA [100] Video Object Detection 5.08 2.64 65.57 48.18
FGFA (FlowNet [86][87] Video Object Detection 0.23 0.12 33.97 46.70
FGFA (PWC-Net [88][87] Video Object Detection 0.20 0.10 32.01 45.54
FGFA (RAFT [89][87] Video Object Detection 0.14 0.07 30.09 43.03
V-A Ablation study on the proposed method
In this part, we conduct ablation studies to evaluate the effectiveness of each module in our proposed method FADE-Net. As shown in Table III, both accuracy and recall are improved when adopting SMRPN. Notably, the recall is boosted by more than 20% (26.34% vs 46.53% ). This indicates that multi-stage cascade refinement enhances the detection accuracy. Additionally, employing a dynamic area-based anchor mining strategy in the initial stage helps avoiding missed detection of small objects. The adoption of MAM resulted in a significant increase of 38.06% in the recall rate. This indicates that incorporating moving object regions as an attention mask in the detection network effectively captures high-speed falling objects. The simultaneous usage of SMRPN and MAM further enhances the detection performance, suggesting that these two modules are complementary. This improvement shows that the ability to detect small objects and capture trajectories of moving objects contributes to the effectiveness of falling object detection around buildings.
V-B Performance in the testing set
@@
To provide a comprehensive benchmark, we conduct extensive experiments on FADE. Various state-of-the-art methods are evaluated, including our method FADE-Net, 5 generic object detection methods, 9 MOD methods and 2 video object detection methods. We use the metrics defined in section III-E for performance evaluation. We specify the implementation details of all algorithms (requirements, hyperparameters, and training details) at https://github.com/Zhengbo-Zhang/FADE.
@@
Our proposed FADE-Net achieves the highest performance across all four metrics (see Table IV). During the training process, the CNN generates feature maps that are significantly smaller than the original image, making it is challenging for generic object detection methods to capture features of small objects. However, the statistical data presented in Table II reveals that the falling objects in our dataset are of small size. Our FADE-Net benefits from multi-stage proposal refinement and area-based anchor mining strategy, enabling better detection of small objects. Furthermore, motion blur caused by the fast motion of falling objects can reduce the recall of image-based detection models. Our FADE-Net focuses on determining the candidate motion object regions in adjacent frames and incorporates motion-attention masks at various stages of the network, providing more accurate detection.
Despite being the best MOD method, MOG2 [42] still exhibits a significant precision gap compared to our proposed FADE-Net. The reason for this is that our FADE-Net learns rich appearance features of falling objects during offline training and effectively extracts these features during the inference process, which greatly aids in falling object detection. In contrast, the MOD method can only update the model online during the inference process and is unable to leverage this valuable prior knowledge. This causes MOG2 to detect some moving objects that are not falling objects, such as clouds and shadows. The detection results of LSBP [99] contain a large number of ghost regions (trailing regions at the tail of the fast moving object which are outside the contour of the fast moving object) [101], leading to its worst precision and F-measure.
MEGA [100] combines the global semantic information with the local localization information, as well as utilizes more key frames in the video for detection. It therefore is very sensitive to detect the falling object and to locate the beginning-ending time of the FODB incident. It achieves the second best recall and TRO performance. Since MEGA is not aimed at small object detection, its network cannot learn the appearance features of the small falling objects well. In addition, since the falling objects move too fast and the motion occupies several video frames, the motion region is easily submerged in the time series. As a result, MEGA gets low precision in the constructed FADE dateset.
Figure 9: Video frames in our FADE dataset and their corresponding optical flow field images estimated by different optical flow methods. (a) Four consecutive video frames with a falling object from the building in our dataset. (b), (c), (d), and (e) The corresponding optical flow images computed by FlowNet 2.0 [102], RAFT [89], PWC-Net [88], and TV-L1 [103]. To see the moving object in the falling object detection task clearly, we enlarge the object in the frame and place it in the lower left corner of the video frame.
@@
V-C Analysis of optical flow based methods in FODB
Optical flow is good at estimating motion information in video by capturing the pixel-level dense motion field. It is widely used in many video tasks, e.g. action recognition [104, 105, 106, 107] and MOD [100], and makes good performance. However, as shown in Table IV, the optical flow based methods FGFA (FlowNet [86][87] (the original method using FlowNet to compute optical flow), FGFA (PWC-Net [88][87], and FGFA (RAFT [89][87], do not obtain the expected performance in falling object detection in our FADE (the F-measure of FGFA (RAFT) is the second worst). To find out the reason, we apply a variety of representative optical flow methods, e.g. FlowNet, PWC-Net, and RAFT to generate the optical flow images in FADE for comparison, and we provide these images in Figure 9. By observing these optical flow images estimated by the deep learning based optical flow algorithms [86, 89, 88], we find that some optical flow images are chaotic, resulting in the background and the falling object being difficult to be distinguished. Besides, the displacement of falling object is inaccurately computed.

### v3

@@ III-C Dataset division and statistics
To provide a more intuitive understanding of our dataset, we present several quantitative statistics, including the dataset division, falling object sizes, the number of instances per object category, and the proportion of object area relative to the image.
Dataset Division. When splitting the dataset, we consider five attributes for each video: video resolution, scene, lighting condition, weather condition, and object category. It is worth noting that the videos in our training, validation, and testing sets collectively cover all labels from the five attributes mentioned above. To better test the generalization of FODB algorithms, the scenes in our training set, validation set, and testing set (except for the scenes captured in rainy days) are non-overlapped.
Dataset Statistics. We provide some statistical information about the constructed dataset. As shown in Figure 5, to make our FADE dataset with various falling object categories, we collect the videos covering falling objects of eight categories. The percentage of falling objects with different weights is shown in Figure 6.
Figure 7: Overview of our proposed method, which is built upon the Faster R-CNN [80] framework with Feature Pyramid Network [81].
“MAM” denotes the Moving attention module. “C” and “A” denote classifier and anchor regressor, respectively. “Conv” and “AdaConv” indicate conventional convolution and the adaptive convolution [82] layers, respectively.
Figure 8: Details of the Moving Attention Module. The Moving Attention Module is designed to fuse the appearance and motion information of objects. To improve the robustness of the appearance features, we apply average pooling and max pooling to the input appearance features before the fusion step.
@@
To facilitate the use of our dataset, we provide details regarding its ethical considerations, licensing terms, and maintenance plan in this section.
Ethics. We have full ownership of these videos supplied in our FADE dataset, and all of the collected videos have been authorized for free use by the corresponding university. We inform the seven volunteers, who help us drop the objects from buildings, in advance that their personal information (e.g. faces and hands) may appear in the video, and we get their permission. We conduct two improvements to solve the ethical concerns in our dataset. 1) We have reviewed all the videos and found that 32 of them contain people. Among these, 21 videos include appearances of volunteers who assisted in the data collection process. For the 21 videos, the volunteers are informed of their appearance and have signed on an informed consent form to allow us to use the videos. The link of this signed informed consent form is: https://fadedataset.github.io/FADE.github.io/ethic.html. For the other 11 videos, we remove them from our FADE dataset. 2) We have obtained authorization from the corresponding university where the video data was captured. All videos in the updated FADE dataset are freely available for research on falling object detection. We upload the corresponding authorization files to our dataset website (https://fadedataset.github.io/FADE.github.io/ethic.html). Notably, the video in our dataset does not contain any personally identifiable information or offensive content, and we assume full responsibility in the event of any issues related to data licensing.
@@ III-E Evaluation metrics
We use precision, recall, and F-measure to evaluate and compare the performance of methods across different tasks, including MOD, generic object detection, and video object detection. Since the object size in our dataset FADE is small, we treat a detection as a true positive if its IoU with the GT is larger than 0.3. The F-measure is defined as:
@@
We set the positive real factor $\beta=1$, when we calculate the F-measure.
@@ IV Proposed Method
Generic object detection methods rely heavily on appearance features to detect objects. However, in the FODB task, falling objects are typically small and often exhibit blurred appearances due to motion blur caused by high-speed movement, making them difficult to detect using appearance features alone. To address this limitation, we propose a method called FADE-Net, built upon Faster R-CNN [80] with FPN [81] (see Figure 7). Specifically, FADE-Net introduces a Moving Attention Module that incorporates motion cues to complement appearance features for improved detection of fast-moving objects. Additionally, a Small-Object Mining Region Proposal Network (SMRPN) is integrated in our FADE-Net to enhance the detection of small objects common in FODB scenarios. The details of MAM and SMRPN are described in the following sections.
IV-A Moving attention module (MAM)
Detecting falling objects around buildings based solely on appearance features is challenging, as these objects are often small and affected by motion blur due to their high speed. To enhance detection accuracy, we leverage the inherent motion characteristics of falling objects by designing a Moving Attention Module that incorporates motion information as a complementary cue to appearance features. Specifically, we first employ a Moving Object Prediction Module to extract a motion mask representing moving objects. This mask is then fed into the proposed Moving Attention Module, where it is fused with appearance features to produce more informative representations for falling object detection. The designs of the Moving Object Prediction Module and the Moving Attention Module are detailed in the following sections.
Moving Object Prediction Module. To obtain motion information as a complementary cue to appearance features for improved falling object detection, we design a Moving Object Prediction Module. Specifically, to balance robustness and real-time performance, we adopt the MOD method MOG2 [45] to generate moving object masks as motion cues. The masks are candidate moving object regions within the current frame.
Moving Attention Module. The generated moving object mask, which serves as motion information, is fed into the Moving Attention Module and fused with appearance features to provide more robust representations for falling object detection. Specifically, in our Moving Attention Module, the appearance feature is represented by the feature map $F$ from the previous layer, with dimensions $H\times W\times C$. To enhance the robustness of the appearance information, we apply both average pooling and max pooling to $F$ before fusion.
We then concatenate the averaged and max-pooled feature maps with the moving object mask and feed them into a convolutional layer followed by a Sigmoid activation to fuse motion and appearance features. The fused output, after passing through the convolutional layer and Sigmoid activation, forms a more robust representation, referred to as the moving attention map, which serves as the output of our Moving Attention Module. The process is defined as:
$M=\sigma(\text{Conv}(\text{Concat}(\text{AvgPool}(F),\text{MaxPool}(F),\text{Mask}))),$ (4)
where $M$ is the resulting moving attention map, and $\sigma$ denotes the Sigmoid function. It is worth noting that our Moving Attention Module is applied at three locations, each positioned after the downsampling stages in our backbone. All Moving Attention Modules share the same moving object mask.
Overall, the proposed Moving Attention Module enables our detection model to effectively integrate appearance and motion information, resulting in more robust representations for detecting falling objects around buildings.
IV-B Small-Object Mining RPN (SMRPN)
After introducing the Moving Attention Module, we now present the Small-Object Mining RPN. While the Moving Attention Module enables the model to leverage both appearance and motion information for more robust detection of small objects, the small size of falling objects may still result in their features being lost during downsampling in the backbone. To address this challenge, we design the Small-Object Mining RPN (SMRPN), which integrates multi-level features and employs dynamic thresholds to ensure that small falling objects can be effectively captured by the model.
Specifically, SMRPN leverages features from all downsampled levels to perform the initial regression, enabling the model to effectively capture information across different object scales. We also adopt an adaptive convolution [82] to align anchor features with target features, allowing for more refined regression and ultimately generating high-quality proposals for small objects. In addition, to enhance the model’s ability to detect typically small falling objects, we introduce an area-based anchor mining strategy with a dynamic threshold that adapts to object size in the first stage:
@@ IV-B Small-Object Mining RPN (SMRPN)
$\begin{split}Threshold=\mathrm{max}\left(0.20,0.15+\alpha\cdot\log{\frac{\sqrt{w\cdot h}}{5}}\right),\end{split}$ (5)
where $\alpha$ is a scale factor, set to 0.2 by default in our task.
TABLE II: Performance of different methods on the testing set of our FADE dataset. The best and second best performances are highlighted in bold and underline, respectively. DLA34 (DCNv2 [84][85] indicates that some of the convolutions in the DLA34 network are replaced by Deformable Convolution v2 (DCNv2). FGFA (FlowNet [86][87], FGFA (PWC-Net [88][87], and FGFA (RAFT [89][87] respectively indicate FGFA based on the optical flow methods of FlowNet, PWC-Net, and RAFT.
Method Type F-measure Precision Recall TRO FPS
FADE-Net (Ours) MOD-based Generic Object Detection 72.08 73.52 70.69 51.77 15.7
Faster R-CNN [80] + FPN [81] Generic Object Detection 35.56 54.55 26.38 32.47 16.7
YOLOv5 [90] Generic Object Detection 33.67 54.77 24.31 34.12 32.8
DLA34 (DCNv2 [84][85] Generic Object Detection 22.57 20.04 25.83 24.69 22.7
DETR [91] Generic Object Detection 29.98 49.47 21.51 26.80 7.9
swin-B [92] Generic Object Detection 36.99 57.37 27.29 36.63 8.3
RT-DETR [93] Generic Object Detection 40.15 59.24 30.36 38.62 28.7
MOG [94] Moving Object Detection 17.96 12.85 29.80 20.91 370.1
MOG2 [45] Moving Object Detection 24.55 19.03 34.57 48.03 331.8
GMG [95] Moving Object Detection 2.71 1.55 10.89 14.71 97.7
Vibe [8] Moving Object Detection 15.26 14.66 15.91 23.85 488.5
CNT [96] Moving Object Detection 12.84 19.09 9.67 17.25 147.3
FMOD [97] Moving Object Detection 2.48 6.73 1.52 16.09 289.7
KNN [98] Moving Object Detection 0.83 0.42 30.48 34.20 188.5
GSOC [99] Moving Object Detection 0.73 0.38 10.12 16.17 146.6
LSBP [100] Moving Object Detection 0.20 0.10 6.73 12.35 158.0
MEGA [101] Video Object Detection 5.20 2.71 65.60 48.23 8.3
FGFA [87] (FlowNet [86]) Video Object Detection 0.26 0.13 34.01 46.72 6.7
FGFA [87] (PWC-Net [88]) Video Object Detection 0.22 0.11 31.95 45.47 10.1
FGFA [87] (RAFT [89]) Video Object Detection 0.18 0.09 30.05 42.97 8.8
@@ V Experiments
In this section, we sequentially present the implementation details of our method, the main results, the ablation study, an analysis of the performance of optical flow-based methods, and visualization results.
V-A Implementation details
Our method is fine-tuned on a pre-trained Faster R-CNN [80] with a ResNet-50 [102] backbone. We use SGD as the optimizer with a learning rate of 0.005, momentum of 0.9, and weight decay of 0.0005. The model is trained for 15 epochs with a batch size of 2 on our dataset. The Moving Attention Module consists of a convolution layer with a 7 $\times$ 7 kernel followed by a Sigmoid function. Its input includes a moving object mask predicted by MOG2 [45] along with its corresponding average pooling map and max pooling map. The kernel size of the convolution layer in our moving attention module is set to $7\times 7$.
V-B Main results
@@
To provide a comprehensive benchmark, we conduct extensive experiments on FADE, evaluating a range of state-of-the-art methods, including our proposed FADE-Net, five generic object detection methods, nine MOD methods, and two video object detection methods. We use the metrics defined in section III-E for performance evaluation. We specify the implementation details of all algorithms (requirements, hyperparameters, and training details) at https://github.com/Zhengbo-Zhang/FADE.
@@
As can be seen in Table II, our proposed FADE-Net achieves the highest performance across all four metrics. During the training process, the CNN generates feature maps that are significantly smaller than the original image, making it is challenging for generic object detection methods to capture features of small objects. However, the falling objects in our dataset are of small size. To address this challenge, our FADE-Net leverages multi-stage proposal refinement and an area-based anchor mining strategy, enabling more effective detection of small objects. Furthermore, motion blur caused by the fast motion of falling objects can reduce the recall of image-based detection models. To solve it, our FADE-Net integrates the proposed Moving Attention Module, which fuses motion information with appearance features to produce more robust representations for falling object detection.
Although MOG2 [45] is the best-performing MOD method, it still shows a substantial precision gap compared to our proposed FADE-Net. This is because FADE-Net effectively extracts and utilizes both appearance and motion features during inference, which significantly enhances falling object detection. In contrast, the MOD method can only update the model online during inference and lacks the ability to exploit various informative features. This leads MOG2 to detect certain moving objects that are not falling objects, such as clouds and shadows. The detection results of LSBP [100] contain numerous ghost regions, such as trailing areas behind fast-moving objects that lie outside their actual contours [103], resulting in the lowest precision and F-measure. MEGA [101] combines global semantic information with local localization cues and leverages more key frames in the video for detection, achieving the second-best recall and TRO performance. However, since MEGA is not designed for small object detection, its network struggles to capture the appearance features of small falling objects. As a result, MEGA yields low precision on the constructed FADE dataset.
The experimental results show that although our method is not the fastest in terms of inference speed, it achieves superior performance owing to the SMRPN and MAM modules, which are specifically designed for FODB. In terms of performance, our method significantly outperforms the second-best approach (F-measure 72.08 vs 40.15).
TABLE III: Performance of different modules in the proposed FADE-Net on the testing set of our FADE dataset. We use a GTX 1080 to evaluate the FPS of the algorithm. The best performances are highlighted in bold.
Faster R-CNN+FPN SMRPN MAM F-measure Precision Recall TRO FPS
$\surd$ 35.62 54.78 26.39 32.50 16.5
$\surd$ $\surd$ 53.42 62.75 46.51 39.19 16.2
$\surd$ $\surd$ 61.72 59.21 64.46 47.25 15.9
$\surd$ $\surd$ $\surd$ 72.08 73.52 70.69 51.77 15.7
@@
V-C Ablation study
In this part, we conduct ablation studies to evaluate the effectiveness of each module in our proposed method FADE-Net. As shown in Table III, both accuracy and recall are improved when adopting SMRPN. Notably, the recall is boosted by more than 20% (26.39% vs 46.51%). This indicates that multi-stage cascade refinement enhances the detection accuracy. Additionally, employing a dynamic area-based anchor mining strategy in the initial stage helps avoiding missed detection of small objects. The adoption of MAM resulted in a significant increase of 38.07% in the recall rate. This indicates that incorporating motion information in our method enables effective capture of high-speed falling objects. The simultaneous usage of SMRPN and MAM further enhances the detection performance, suggesting that these two modules are complementary. Such improvement also shows that the ability to detect small objects and capture trajectories of moving objects contributes to the effectiveness of falling object detection around buildings. Besides, the experimental results show that introducing the SMRPN and MAM modules has minimal impact on the inference speed of our method. This is because the MAM module is built upon an efficient GPU-accelerated MOG algorithm, while the SMRPN functions primarily as a training-time strategy that employs dynamic thresholds to generate proposals tailored for small object detection.
V-D Analysis of optical flow based methods in FODB
Optical flow is good at estimating motion information in video by capturing the pixel-level dense motion field. It is widely used in many video tasks, e.g. action recognition [104, 105, 106, 107] and MOD [101], and makes good performance. However, as shown in Table II, the optical flow based methods FGFA (FlowNet [86][87] (the original method using FlowNet to compute optical flow), FGFA (PWC-Net [88][87], and FGFA (RAFT [89][87], do not obtain the expected performance in falling object detection in our FADE (the F-measure of FGFA (RAFT) is the second worst).